# Mean field equation and Green’s function on the Tate curve

### 16 Mar 2026 v1

@@ 3.1 Structure of the solution
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{v_{p}(z)=l}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$ (3.9)
$+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}-(p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0$
Fixing $b_{0},b_{1},\dots,b_{d-3}$ and varying $b_{d-2}$, we see that $u(x)$ satisfies (LABEL:1_5). So when $\rho p^{2d}\leq p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-2})$ is constant for all $b_{d-2}$, denote it by $u_{l}(b_{0},b_{1},\dots,b_{d-3})$. Similarly, if we fix $b_{0},b_{1},\dots,b_{d-4}$ and when $\rho p^{2d}\leq p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-3})$ is constant for all $b_{d-3}$. Proceeding recursively, we obtain
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{z}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},\dots,a_{s},a_{s+1})$ (3.10)
$+\rho p^{d}e^{u_{l}(b_{0},\dots,b_{s},b_{s+1})}-(p^{d+1+s}+p^{2d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},\dots,b_{s},b_{s+1})=0$
@@
If we fix $b_{i}=c_{i}$, $0\leq i\leq s$, $b_{s+1}\neq c_{s+1}$ and suppose $u_{l}(c_{0},\dots,c_{s},b_{s+1},\dots,b_{d-1})$ is independent of $b_{s+2},\dots,b_{d-1}$, denote it by $u_{l}(c_{0},\dots,c_{s},b_{s+1})$. Next, we prove that $u_{l}(c_{0},\dots,c_{s},b_{s+1})$ is independent of $b_{s+1}$: From (3.6) and (LABEL:coefficient_of_u(x)), we can get
$p^{d}Du(x)=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$ (3.11)
@@
$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq s\\ a_{s+1}\neq b_{s+1}\\ a_{s+2}\dots a_{d-1}\end{subarray}}p^{2(s+1)}u(z)$
$+\sum_{j=s+2}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
we compute
@@
$=$ $p^{2(s+1)}(p^{d-s-2}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
$+(p-1)p^{d-s-3}u_{l}(c_{0},\dots,c_{s+1})+\dots+(p-1)u_{l}(c_{0},\dots,c_{d-2})+u(y))$
@@
$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
$+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)-(p^{d+s+1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
$=$ $\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
$+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)-(p^{d+s+1}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
so when $p^{2d}\rho\leq p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}$ depend only on $c_{0},c_{1},\dots,c_{s}$, $\forall s=0,1,\dots,d-2$. Then we get the result by induction. In conclusion, when $v_{p}(x)=l=v_{p}(y)$, $p^{2d}\rho\leq p^{d+1+s}+p^{2d-2}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(c_{0},\dots,c_{s})$ is constant. Then when
$\rho$ $\leq p^{-2d}\text{min}_{\begin{subarray}{c}-1\leq s\leq d-2\\ 0\leq l\leq m-1\end{subarray}}\{p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}\}$
@@
$=p^{-d}-p^{-d-m}$
$u(x)$ is radial symmetric. By Lemma 3.4 and Lemma 3.5, we can see that $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$, and if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ with $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$, which means that $u(x)$ is radial symmetric. ∎
From our numerical results, we can also see that the radial symmetric solutions have monotonicity. $u(x)$ increases with the distance from the singular point $y$.
@@
Assume $\rho\leq p^{-d}-p^{-d-m}$ and $v_{p}(x)=l=v_{p}(y)$, then $u_{l}(c_{0},\dots,c_{d-1})<u_{l}(c_{0},\dots,c_{d-2})<\dots<u_{l}(c_{0})<u_{l}$
@@
Let $v_{p}(z)=k$, by Theorem 3.6, when $\rho\leq p^{-d}-p^{-d-m}$, $u(x)$ is radial symmetric, so the equation variables can be reduced from $m(p-1)p^{d-1}$ to $p+m$. Then denote these variables by $u_{0},\dots,u_{l-1},u_{l},u_{l}(c_{0}),\dots,u_{l}(c_{0},\dots,c_{d-1}),u_{l+1},\dots,u_{m-1}$.
When $x\neq y$, equation (3.4) becomes
$\sum_{k=0}^{l-1}p^{k-l}u(z)+\sum_{k=l+1}^{m-1}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}+p^{d}\rho e^{u_{l}(x)}=0$ (3.13)
When $b_{i}=c_{i}$ for $0\leq i\leq s$, then from (LABEL:Du_3), we have
@@
$+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
$=$ $\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
@@
Suppose $u_{l}(c_{0},\dots,c_{s+1})>\dots>u_{l}(c_{0},\dots,c_{d-1})$, now we want to compare $u_{l}(c_{0},\dots,c_{s+1})$ with $u_{l}(c_{0},\dots,c_{s})$. First, by (LABEL:eq2_1) and (LABEL:eq2_2), they satisfy the following equations:
$(p-1)p^{d-1}(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k})+\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.15)
$+p^{2(s+1)}(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
@@
$(p-1)p^{d-1}(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k})+\sum_{j=0}^{s+1}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.16)
$+p^{2(s+2)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1}))$
$-(2p^{s+2}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})+p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}=0$
subtracting (LABEL:eq2_4) from (LABEL:eq2_3) gives
@@
$+(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})+(p-1-\delta_{s+1,0})p^{d+s}u_{l}(c_{0},\dots,c_{s})$
@@
$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
@@
$(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}+1)$
$=$ $(p^{2}-1)p^{2(s+1)}((p-1)p^{d-2}\frac{p^{-s-1}-p^{-d+2}}{p-1}+1)$
$=$ $p^{d-1}(p^{s+2}-p^{s})$
@@
$+p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s})$
since $\rho\leq p^{-d}-p^{-d-m}$, then by the monotonicity of $f_{s}(x)=\rho e^{x}-p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})x$, we have $u_{l}(c_{0},\dots,c_{s+1})<u_{l}(c_{0},\dots,c_{s})$.
@@
$=$ $(p^{2}-1)p^{2(s+1)}(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}((u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{j}))$
$+(u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{d-1}))$
@@
This theorem provides a sharper estimate of the upper bound for the solution. We now refine Lemma 3.2:
@@
$u(i)<\begin{cases}\text{ln}\frac{p}{p-2},\ &p>2\\ 4,&p=2\end{cases}$.
@@
thus $M<\text{ln}\frac{p}{p-2}$.
When $p=2$, there are at least $2^{d-2}$ such $M$, so $M<4$. ∎
Using this improved bound, we refine Theorem (3.6) and Theorem (3.7):

### 08 Apr 2026 v2

@@ 3.1 Structure of the solution
$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=l\end{subarray}}u(z)+\sum_{j=0}^{d-3}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},a_{1},\dots,a_{d-2})$ (3.11)
$+\rho p^{d}e^{u_{l}(b_{0},b_{1},\dots,b_{d-2})}-(p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},b_{1},\dots,b_{d-2})=0.$
Fixing $b_{0},b_{1},\dots,b_{d-3}$ and varying $b_{d-2}$, we see that $u(x)$ satisfies (LABEL:eq:equation_for_vpx_neq_vpy_d-2). So when $\rho p^{2d}\leq p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-2})$ is constant for all $b_{d-2}$, denote it by $u_{l}(b_{0},b_{1},\dots,b_{d-3})$. Similarly, if we fix $b_{0},b_{1},\dots,b_{d-4}$ and when $\rho p^{2d}\leq p^{2d-2}+p^{2d-3}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(b_{0},b_{1},\dots,b_{d-3})$ is constant for all $b_{d-3}$. Proceeding recursively, we obtain
$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{\begin{subarray}{c}z\\ v_{p}(z)=l\end{subarray}}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j\\ a_{j+1}\dots a_{d-1}\end{subarray}}(p^{2(j+1)}-p^{2j})u_{l}(a_{0},\dots,a_{s},a_{s+1})$
$+\rho p^{d}e^{u_{l}(b_{0},\dots,b_{s},b_{s+1})}-(p^{d+1+s}+p^{2d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(b_{0},\dots,b_{s},b_{s+1})=0,$
@@
If we fix $b_{i}=c_{i}$, $0\leq i\leq s$, $b_{s+1}\neq c_{s+1}$ and suppose $u_{l}(c_{0},\dots,c_{s},b_{s+1},\dots,b_{d-1})$ is independent of $b_{s+2},\dots,b_{d-1}$, denote it by $u_{l}(c_{0},\dots,c_{s},b_{s+1})$. Next, we prove that $u_{l}(c_{0},\dots,c_{s},b_{s+1})$ is independent of $b_{s+1}$: From (3.8) and (LABEL:eq:coefficient_of_u(x)), we can get
$p^{d}Du(x)=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$ (3.12)
@@
$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)+\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq s\\ a_{s+1}\neq b_{s+1}\\ a_{s+2}\dots a_{d-1}\end{subarray}}p^{2(s+1)}u(z)$
$+\sum_{j=s+2}^{d-1}\sum_{\begin{subarray}{c}a_{i}=b_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq b_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1}).$
We first compute
@@
$=$ $p^{2(s+1)}\Biggl(p^{d-s-2}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})$
$+(p-1)p^{d-s-3}u_{l}(c_{0},\dots,c_{s+1})+\dots+(p-1)u_{l}(c_{0},\dots,c_{d-2})+u(y)\Biggr)$
@@
$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
$+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq b_{s+1}\\ a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
$-(p^{d+s+1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})$
$=$ $\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
$+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
$-(p^{d+s+1}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1}).$
Then equation (LABEL:eq:equation_for_x_neq_y) becomes
$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{s}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$
$+p^{d+s}\sum_{\begin{subarray}{c}a_{s+1}\neq c_{s+1}\end{subarray}}u_{l}(c_{0},\dots,c_{s},a_{s+1})+\sum_{j=s+1}^{d-2}(p-1)p^{d+2s-j}u_{l}(c_{0},\dots,c_{j})+p^{2(s+1)}u(y)$
$-(p^{d+s+1}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}(c_{0},\dots,c_{s},b_{s+1})+\rho p^{d}e^{u_{l}(c_{0},\dots,c_{s},b_{s+1})}=0.$
So when $p^{2d}\rho\leq p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}$ depend only on $c_{0},c_{1},\dots,c_{s}$, $\forall s=0,1,\dots,d-2$. Then we get the result by induction. In conclusion, when $v_{p}(x)=l=v_{p}(y)$, $p^{2d}\rho\leq p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}$, $u_{l}(c_{0},\dots,c_{s})$ is constant. Then when
$\rho$ $\leq p^{-2d}\min_{\begin{subarray}{c}-1\leq s\leq d-2\\ 0\leq l\leq m-1\end{subarray}}\{p^{d+1+s}+p^{d+s}-(p^{-l}+p^{l+1-m})p^{d-1}\}$
@@
$=p^{-d}-p^{-d-m},$
$u(x)$ is radially symmetric. By Lemma 3.4 and Lemma 3.6, we can see that $u_{y}(x)=u(\text{d}(x,y),|x|_{p},|y|_{p})$, and if $\text{d}(x_{1},y_{1})=\text{d}(x_{2},y_{2})$ with $v_{p}(x_{1})+v_{p}(x_{2})=v_{p}(y_{1})+v_{p}(y_{2})=m-1$, then $u_{y_{1}}(x_{1})=u_{y_{2}}(x_{2})$. ∎
By observing these radially symmetric solutions from numerical experiments, we find that the solution has monotonicity. $u(x)$ increases with the distance from the singular point $y$.
@@
Assume $\rho\leq p^{-d}-p^{-d-m}$ and $v_{p}(x)=l=v_{p}(y)$, then the radially symmetric solution satisfies $u_{l}(c_{0},\dots,c_{d-1})<u_{l}(c_{0},\dots,c_{d-2})<\dots<u_{l}(c_{0})<u_{l}$
@@
From Theorem 3.7, when $\rho\leq p^{-d}-p^{-d-m}$, the solution is radially symmetric, so the equation variables can be reduced from $m(p-1)p^{d-1}$ to $p+m$. Then denote these variables by $u_{0},\dots,u_{l-1},u_{l},u_{l}(c_{0}),\dots,u_{l}(c_{0},\dots,c_{d-1}),u_{l+1},\dots,u_{m-1}$.
When $x\neq y$, by letting $v_{p}(z)=k$, equation (3.4) becomes
$\sum_{\begin{subarray}{c}z\\ 0\leq k\leq l-1\end{subarray}}p^{k-l}u(z)+\sum_{\begin{subarray}{c}z\\ l+1\leq k\leq m-1\end{subarray}}p^{l-k}u(z)+\sum_{j=0}^{d-1}\sum_{\begin{subarray}{c}a_{i}=c_{i}\\ 0\leq i\leq j-1\\ a_{j}\neq c_{j}\\ a_{j+1}\dots a_{d-1}\end{subarray}}p^{2j}u(z)$ (3.14)
$-(p^{2d-1}-(p^{-l}+p^{l+1-m})p^{d-1})u_{l}+p^{d}\rho e^{u_{l}(x)}=0.$
When $b_{i}=c_{i}$ for $0\leq i\leq s$, then from (LABEL:eq:compute_p^dDu_index_simplified), we have
@@
$+p^{2(s+1)}\left(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
$-(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})$
$=$ $\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})+p^{2(s+1)}\left(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
@@
Suppose $u_{l}(c_{0},\dots,c_{s+1})>\dots>u_{l}(c_{0},\dots,c_{d-1})$, now we want to compare $u_{l}(c_{0},\dots,c_{s+1})$ with $u_{l}(c_{0},\dots,c_{s})$. First, by (LABEL:eq:equation_radial_symmetric_x_neq_y) and (LABEL:eq:equation_radial_symmetric_x_neq_y_crutial_part), they satisfy the following equations:
$(p-1)p^{d-1}\left(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k}\right)+\sum_{j=0}^{s}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.16)
$+p^{2(s+1)}\left(\sum_{j=s+1}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
@@
and
$(p-1)p^{d-1}\left(\sum_{k=0}^{l-1}p^{k-l}u_{k}+\sum_{k=l+1}^{m-1}p^{l-k}u_{k}\right)+\sum_{j=0}^{s+1}(p-1-\delta_{j0})p^{d-1+j}u_{l}(c_{0},\dots,c_{j-1})$ (3.17)
$+p^{2(s+2)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
$-(2p^{s+2}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s+1})+p^{d}\rho e^{u_{l}(c_{0},\dots,c_{s+1})}=0.$
Subtracting (LABEL:eq:equation_radial_symmetric_up_to_s+1) from (LABEL:eq:equation_radial_symmetric_up_to_s) gives
@@
$+(2p^{s+1}+\delta_{s,-1}-p^{-l}-p^{l+1-m})p^{d-1}u_{l}(c_{0},\dots,c_{s})+(p-1-\delta_{s+1,0})p^{d+s}u_{l}(c_{0},\dots,c_{s}),$
@@
$+(p^{2}-1)p^{2(s+1)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}u_{l}(c_{0},\dots,c_{j})+u_{l}(c_{0},\dots,c_{d-1})\right)$
$+p^{d-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s}).$
@@
$(p^{2}-1)p^{2(s+1)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}+1\right)$
$=$ $(p^{2}-1)p^{2(s+1)}\left((p-1)p^{d-2}\frac{p^{-s-1}-p^{-d+2}}{p-1}+1\right)$
$=$ $p^{d-1}(p^{s+2}-p^{s}),$
@@
$+p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})u_{l}(c_{0},\dots,c_{s}).$
Then by the monotonicity of $f_{s}(x)=\rho e^{x}-p^{-1}(p^{s+2}+p^{s+1}-p^{-l}-p^{l+1-m})x$, if $\rho\leq p^{-d}-p^{-d-m}$, we have $u_{l}(c_{0},\dots,c_{s+1})<u_{l}(c_{0},\dots,c_{s})$.
@@
$=$ $(p^{2}-1)p^{2(s+1)}\left(\sum_{j=s+2}^{d-2}(p-1)p^{d-2-j}((u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{j})\right)$
$+(u_{l}(c_{0},\dots,c_{s+1})-u_{l}(c_{0},\dots,c_{d-1})),$
@@
Remark 3.10.
There always exists radially symmetric solutions as we can view the $p+d$ points as a finite weighted graph. Then from the results in [13] we obtain the existence of the solution, regardless of the value of $\rho$.
Remark 3.11.
The upper bound of $\rho$ that ensures $u$ is radially symmetric and monotone is highly dependent on the upper bound of $u$. And Lemma 3.2 yields only a crude estimate for the upper bound of $u$.
From Theorem 3.8, we can provides a sharper estimate of the upper bound for the solution. We now refine Lemma 3.2:
@@
If $\rho\leq p^{-d}-p^{-d-m}$, then the radially symmetric solution $u$ satisfies
$u(i)<\begin{cases}\ln\left(\frac{p}{p-2}\right),\ &p>2;\\ \ln 4,&p=2.\end{cases}$
@@
thus $M<\ln\left(\frac{p}{p-2}\right)$.
When $p=2$, there are at least $2^{d-2}$ such $M$, so $M<\ln 4$. ∎
Using this improved bound, we can refine Theorem (3.7) and Theorem (3.8) by method of continuity: